OpenStudy (experimentx):

Evaluate: \[ \int_0^\infty\frac{1-\cos x}{x^2(x^2+1)}\,dx \]

OpenStudy (anonymous):

integral calculus?

OpenStudy (experimentx):

yep ... very nice problem!!

OpenStudy (anonymous):

haha .. it's hard . hmm, wat grade r u in?

OpenStudy (experimentx):

university!!

OpenStudy (anonymous):

sorry .. wat university?

OpenStudy (experimentx):

after school you go to university :D

OpenStudy (anonymous):

haha . FUNNY!

hartnn (hartnn):

@experimentX does the solution involve taking Laplace Transform and in the end putting s=0 ?

OpenStudy (experimentx):

yeah i found the solution somewhat like that ...

OpenStudy (experimentx):

the other is to make use of Jordan' Lemma

hartnn (hartnn):

even taking laplace of that function is difficult....

OpenStudy (experimentx):

currently I'm doing complex integration ... probably my last problem. moving to another chapter after this.

OpenStudy (anonymous):

\[I(a)=\int_0^\infty\frac{1-\cos ax}{x^2(x^2+1)}\,dx\]\[I''(a)=\int_0^\infty\frac{\cos ax}{1+x^2}\,dx\]

OpenStudy (anonymous):

by jordan\[I''(a)=\int_0^\infty\frac{\cos ax}{1+x^2}\,dx=\frac{\pi}{2e^a}\]?

OpenStudy (anonymous):

santosh?

OpenStudy (experimentx):

the latter part seems simple .. since the residue is only ... +i

OpenStudy (anonymous):

yeah ... is that right?im not sure

OpenStudy (experimentx):

let me evaluate it .. since the function is even ...